# Giulia Gaggero and Elisa Gorla

### v2

@@
The complexity of solving a random polynomial systemthanks: This research was funded by armasuisse through grant no. CYD-C-2020010.
Giulia Gaggero and Elisa Gorla
@@ 1 Introduction
Randomness plays a fundamental role within cryptography. For example, it plays a pivotal role within key generation and cryptographic algorithms are subject to randomness tests. In multivariate cryptography, the public key consists of a multivariate polynomial system and the goal of the attacker is to find a solution of the system. In this context, one typically wishes for the system be as close as possible to, or at least appear as, a random system. A random system is expected to be hard to solve, since the Multivariate Quadratic Problem is not only NP-complete, but also known to be hard to solve on average for a wide range of parameters. From this point of view, e.g., a digital signature scheme whose public keys are sufficiently random is expected to be secure.
In this paper, we discuss what it means for a polynomial system to be random and how hard it is to solve a random polynomial system. In Definition 7 we propose a mathematical formulation for the concept of random system. The definition of randomness that we propose, which we call algebraic randomness, is broad enough to include a vast majority of the systems which are of interest in cryptography. We then specify our definition further in Definition 12. One advantage of this definition is that the property of being random according to Definition 12 can be computationally tested, at least in principle.
In Theorem 22, Corollary 26, Corollary 28, Theorem 34, and Proposition 32, we prove upper bounds for the degree of regularity and the solving degree of an algebraically random polynomial system, depending on parameters of the system such as the number of equations, the number of variables, and the degrees of the equations. The usefulness of our bounds is twofold: On the one side, our bounds can be used to directly produce bounds on the complexity of computing a Gröbner basis, hence of solving, many systems which are of interest in cryptography. Bounds on the complexity produced in this way have the advantage of being widely applicable and the disadvantage of not always being close to the actual complexity for each system to which they apply. On the other side, our bounds give us an idea of what security one can hope to achieve for a system with given parameters. Therefore, our bounds can be used as a point of comparison for the optimality of a given public key, in the following sense. Say that, in order to forge a signature produced with a given multivariate digital signature scheme, one has to find a solution of a system of $m$ equations of degree $D$ in $n$ variables. Suppose that such a system is algebraically random. Our results provide an upper bound $B$ for the degree of regularity or the solving degree of such a system. Say that one can compute or estimate by a different method the degree of regularity or the solving degree of the specific system and suppose that this turns out to be $C$. Clearly, it must always be that $C\leq B$. However, how far $C$ is from $B$ gives us a measure of how close to optimal the digital signature scheme is, for the given choice of parameters. In fact, since our bounds on the degree of regularity are sharp, there are systems of $m$ equations of degree $D$ in $n$ variables whose degree of regularity is exactly $B$. In other words, if $B$ and $C$ are close, then there is not much space for improvement, since any system with the same parameters as our public key can have degree of regularity or solving degree at most $B$. If on the contrary $B$ and $C$ are far apart, then potentially there is a lot of space for finding a more robust system with the same parameters, since a system with those parameters can have degree of regularity or solving degree up to $B$.
@@
The paper is structured as follows. Section 2 contains some useful preliminaries. In Section 3 we discuss the concept of randomness and propose two definitions of randomness for a polynomial system, in Definition 7 and Definition 12. Throughout the paper, we refer to our notion of randomness as algebraic randomness. We briefly review cryptographic semiregular sequences and semiregular sequences and show that the latter are algebraically random. In Section 4, after recalling some results from commutative algebra, we prove some bounds on the degree of regularity and the solving degree of an algebraically random system consisting of polynomials the same degree. The bounds appear in Theorem 22, Corollary 26, Corollary 28, Theorem 34, and Proposition 32. Finally, in Section 5 we show how to apply our results to check that the systems associated to GeMSS and Rainbow are not algebraically random.
@@ 2 Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of cardinality $q$ and $R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be the polynomial ring over $\mathbb{F}_{q}$ in $n$ variables. We start by recalling how a typical multivariate one-way function is constructed. Let $f_{1},\ldots,f_{m}\in R$, and consider the evaluation map
@@
To hide the structure of $\mathcal{F}$, we compose it with two random invertible linear maps $S:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n}$ and $T:\mathbb{F}_{q}^{m}\rightarrow\mathbb{F}_{q}^{m}$. We obtain $\mathcal{P}=T\circ\mathcal{F}\circ S$, a set of $m$ polynomials $p_{1},\ldots,p_{m}$ in $n$ variables over $\mathbb{F}_{q}$. The public key of the multivariate scheme is $\mathcal{P}=(p_{1},\ldots,p_{m})$ and the private key is $\{\mathcal{F},S,T\}$. The trapdoor consists of constructing $\mathcal{F}$ such that $\mathcal{F}^{-1}$ is efficiently computable. Notice that $\mathcal{P}$ should be hard to invert without the knowledge of $S,T$, in particular it should be hard to recover the structure of $\mathcal{F}$ from $\mathcal{P}$.
@@
Finding inverse images with respect to $\mathcal{P}$ corresponds to computing solutions of multivariate polynomial systems. It is well-known that this problem can be solved by computing a Gröbner basis of the system, we refer the interested reader to [13] for more detail. The first algorithm for computing Gröbner bases appeared in the doctoral thesis of Buchberger [11]. Modern algorithms for computing Gröbner bases are based on liner algebra and are more efficient than Buchberger’s. Examples of linear-algebra-based algorithms are $F_{4}$ [20], $F_{5}$ [21], the XL Algorithm [18], MutantXL [12], and their variants. In all of these systems, one computes the reduced row echelon form of the Macaulay matrix associated to the polynomial equations in a given degree, for one or more degrees.
@@
We now describe the main object of linear-algebra-based algorithms [6, Section 1.2]. Fix a term order on $R$ and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subseteq R$ be a system of homogeneous polynomials. The columns of the homogeneous Macaulay matrix $M_{d}$ of $\mathcal{F}$ are labelled by the monomials of $R_{d}$ and arranged in decreasing order. The rows of $M_{d}$ are labelled by polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})=d$. The entry $(i,j)$ of $M_{d}$ is the coefficient of the monomial of column $j$ within the polynomial corresponding to the $i$-th row.
@@
Let now $f_{1},\dots,f_{m}$ be arbitrary (not necessarily homogeneous) polynomials. The columns of the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}$ are labelled by the monomials of $R$ of degree $\leq d$, arranged in decreasing order. The rows of $M_{\leq d}$ correspond to polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})\leq d$. The entries of $M_{\leq d}$ are defined as in the homogeneous case. The rationale behind the use of homogeneous Macaulay matrices for homogeneous systems is that, for a homogeneous system, the Macaulay matrix $M_{\leq d}$ is a block matrix with blocks $M_{d},\ldots,M_{0}$ as blocks.
@@
One computes the reduced row echelon form of the Macaulay matrix, or of its homogeneous version, in one or more degrees. For large enough degree, this produces a reduced Gröbner basis with respect to the chosen order. Some algorithms, as e.g. MutantXL, use a variation called mutant strategy in the non-homogeneous case: If the reduction of the Macaulay matrix $M_{\leq d}$ produces new polynomials $g_{1},\ldots,g_{\ell}$ of degree strictly smaller than $d$, one appends to the reduction of $M_{\leq d}$ the polynomials $m_{i,j}g_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}g_{j})\leq d$, then computes the reduced row echelon form again. Throughout the paper, we refer to the algorithms that employ the mutant strategy as mutant algorithms and to the others as standard algorithms. For a description of the basic version of these algorithms, we refer to [13, Section 3.1] and [14, Section 1].
@@
The computational complexity of computing the reduced row echelon form of the Macaulay matrices $M_{\leq d}$ and $M_{d}$ depends on their size, and hence on the degree $d$. This motivates the next definition.
@@
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ and let $\tau$ be a term order on $R$. The solving degree of $\mathcal{F}$ with respect to $\tau$ is the least degree $d$ such that Gaussian elimination on the Macaulay matrix $M_{\leq d}$ produces a $\tau$-Gröbner basis of $\mathcal{F}$. We denote by $\operatorname{\mathrm{solv.deg}}_{\tau}^{s}(\mathcal{F})$ the solving degree of $\mathcal{F}$ with respect to a standard algorithm and by $\operatorname{\mathrm{solv.deg}}_{\tau}^{m}(\mathcal{F})$ the solving degree of $\mathcal{F}$ with respect to a mutant algorithm. When $\tau$ is the degree reverse lexicographic order, we omit the subscript $\tau$.
In general, the solving degree is not invariant under coordinate change. Moreover, it may depend on the algorithm used to perform the Gröbner basis computation. In particular, for mutant algorithms it may be smaller than for standard ones. Finally, the solving degree depends on the choice of a term order on $R$.
The complexity of linear-algebra-based algorithms is dominated by the cost of computing a degree reverse lexicographic Gröbner basis of the system, see [13, Sections 2 and 3] for more detail. Therefore, an upper bound on the solving degree with respect to the reverse lexicographic order yields an upper bound on the complexity of computing a lexicographic Gröbner basis, hence on the complexity of solving the polynomial system.
Let $I$ be a homogeneous ideal of $R$. For an integer $d\geq 0$, we denote by $I_{d}=I\cap R_{d}$ the $\mathbb{F}$-vector space of homogeneous polynomials of degree $d$ in $I$. For $g\in R$ a polynomial, we denote by $g^{\operatorname{\mathrm{top}}}$ the homogeneous part of $g$ of largest degree. E.g., if $g=x^{3}+2xy^{2}-y+1\in\mathbb{F}[x,y]$, then $g^{\operatorname{\mathrm{top}}}=x^{3}+2xy^{2}$. For a polynomial system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$, we denote by $\mathcal{F}^{\operatorname{\mathrm{top}}}\subseteq R$ the homogeneous system $\{f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}\}$. Up to doing Gaussian elimination in a matrix whose rows correspond to $f_{1},\ldots,f_{m}$, we may suppose that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent. We will assume this throughout the paper.
@@
In the cryptographic literature, the degree of regularity is often used as a proxy for the solving degree. This is the case, e.g., in the specification documents of GeMSS [15]. However, this does not always produce reliable estimates. In fact, there are examples in which the gap between the degree of regularity and the solving degree is large, see e.g. [10, Examples 3.2 and 3.3]. A recent result by Semaev and Tenti [27, 28] however shows that, under suitable assumptions, the solving degree of a system with respect to a standard algorithm is at most twice the degree of regularity. Thanks to this result, an upper bound for the degree of regularity yields a proven upper bound for the solving degree.
Theorem 3 ([28, Corollary 3.67] and [27, Theorem 2.1]).
Let $\mathbb{F}=\mathbb{F}_{q}$. Let =$\mathcal{F}=\{f_{1},\ldots,f_{m},x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\subseteq R$ be a polynomial system. If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then for standard algorithms
$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq 2d_{\operatorname{reg}}(\mathcal{F})-2.$
Notice that, in almost all systems of cryptographic interest, the field size and the degrees of the polynomials are relatively small. Therefore, one expects that Theorem 3 applies to such systems.
@@
Theorem 4 ([26, Proposition 3.10]).
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system, then for mutant algorithms
$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq\max\{d_{\operatorname{reg}}(\mathcal{F})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}.$
Another algebraic invariant connected to the solving degree of a polynomial system is the Castelnuovo-Mumford regularity. We refer the reader to [13, Section 3.4] and [14] for its definition and a discussion on its relation with the other invariants.
Remark 5.
If $\mathbb{F}=\mathbb{F}_{q}$ and the system $\mathcal{F}$ contains the field equations $x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}$, then $\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}$ for $d\gg 0$, therefore the degree of regularity is finite.
Remark 6.
If $\left(\mathcal{F}^{\mathrm{top}}\right)_{d}=R_{d}$ for $d\gg 0$, then
@@
Therefore:
  • By Theorem 4
    $\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq\max\{\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}$
    for mutant algorithms.
  • If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then by Theorem 3
    $\operatorname{\mathrm{solv.deg}}(\mathcal{F})\leq 2\operatorname{reg}(\mathcal{F}^{\operatorname{\mathrm{top}}})-2$
    for standard algorithms.
3 Random polynomial systems
Consider a system of $m$ equations of degrees $d_{1},\ldots,d_{m}$ in $n$ variables. Over a finite field, one often defines a random system as a polynomial system whose coefficients are chosen uniformly at random in the given field. In this paper, we propose a different definition which still captures the intuitive idea of randomness, while allowing us to estimate the degree of regularity of a random system.
Over an infinite field, one may use the concept of genericity from algebraic geometry to define randomness. More precisely, fix a nonempty Zarisky-open subset of $\mathbb{P}^{\binom{n+d_{1}}{n}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}}{n}-1}$, where $\mathbb{P}^{t}$ denotes $t$-dimensional projective space. Define a random system as an element of that open set. This makes sense, since every nonempty Zarisky-open set is dense, hence a system of $m$ equations of degrees $d_{1},\ldots,d_{m}$ in $n$ variables whose coefficients are chosen uniformly at random is generic with high probability according to this definition. The problem of extending this definition to a finite field is that, over a finite field, a nonempty Zarisky-open set is no longer dense, so the connection with our intuitive idea of randomness is lost. Nevertheless, for a finite field $\mathbb{F}_{q}$, we may define randomness using a Zariski-dense open set defined over the algebraic closure $\overline{\mathbb{F}_{q}}$. While it is not necessarily the case that almost every polynomial system of given degrees with coefficients in $\mathbb{F}_{q}$ is random, this is the case whenever $q$ is large enough, or if we consider a finite extension of $\mathbb{F}_{q}$ of large enough cardinality.

### v3

@@
Abstract
Many systems of interest in cryptography consist of equations of the same degree. Under the assumption that the degree of regularity is finite, we prove upper bounds on the degree of regularity of a system of equations of the same degree, with or without adding the field equations to the system. The bounds translate into upper bounds on the solving degree of the systems, and hence on the complexity of solving them via Gröbner bases methods. Our bounds depend on the number of equations in the system, the number of variables, and the degree of the equations.
@@ 1 Introduction
Polynomial systems play an important role within cryptography, as several cryptographic schemes rely on the difficulty of solving multivariate polynomial equations over finite fields, especially in the post‑quantum setting. Most notably, multivariate cryptography relies on the difficulty of solving multivariate polynomial equations over finite fields as its core security assumption. The public key of a multivariate system is a polynomial system, which is easy to evaluate, but is believed to be infeasible to solve without secret structural information, giving rise to candidate quantum‑resistant encryption and signature schemes. Beyond multivariate schemes, computational problems can often be modeled via polynomial systems, whose solutions provide a solution to the original problem. MinRank [15] is an example of such a problem, and the most efficient way of solving it currently consists of modeling it via a polynomial system [5, 2, 4]. In algebraic cryptanalysis, block ciphers, hash functions, and other primitives are sometimes translated into large systems of polynomial equations, whose solutions allow an attacker to recover keys or find collisions, see e.g. [33, 40, 1, 22, 24, 34, 18].
The security of many cryptographic schemes, therefore, relies on the complexity of solving the related polynomial system. The degree of regularity of a polynomial system provides an upper bound on the complexity of solving the system using Gröbner bases [38], and hence an upper bound on the security of the corresponding scheme. However, often the degree of regularity is hard to compute or even estimate. In practice, one cannot compute it directly for systems of cryptographic size, as this would require a Gröbner basis computation for values of the parameters for which it is considered infeasible. Depending on the system in question, the degree of regularity and other related invariants may also be hard to estimate theoretically. Due to these difficulties, cryptographers often rely on heuristic estimates. A widespread approach consists of assuming that a system is semiregular (see e.g. [19]), which provides an explicit formula for its degree of regularity [7].
In this paper, we consider systems for which the degree of regularity is finite. These are exactly the systems for which the approach via the degree of regularity provides a non-trivial complexity bound. For polynomials in $n$ variables, the degree of regularity is finite if and only if the ideal generated by the top-degree part of the system contains a regular sequence of $n$ polynomials. For a system of polynomials of the same degree $D$, the degree of regularity is finite if and only if the regular sequence may be found in degree $D$. For such systems, in Theorem 23 we prove a sharp upper bound for the degree of regularity in terms of the number of equations, the number of variables, and the degree of the equations. In Theorem 33 we prove a similar result for a system of equations to which we add the field equations.
Our bounds on the degree of regularity readily produce bounds on the complexity of computing a Gröbner basis, hence of solving, many systems that are of interest in cryptography. Bounds on the complexity produced in this way have the advantage of being widely applicable and the disadvantage of not always being close to the actual complexity of solving the systems to which they apply. In particular, the complexity estimates they produce are less sharp than those produced using the heuristic assumption that the system is semiregular. However, our bounds are proven and do not rely on any heuristic assumption. In particular, they apply to the situations where no semiregular sequence exists. This is the case for certain choices of the parameters and over a given field. On the other hand, since they are sharp, our bounds give us an idea of what security one can hope to achieve for a system with given parameters.
@@
The paper is structured as follows. Section 2 contains some useful preliminaries. In Section 3 we briefly review cryptographic semiregular sequences and semiregular sequences, and we introduce the family of polynomial systems that are regular in a given degree $D$. These are the systems that we will focus on in this paper. In Section 4, after recalling some results from commutative algebra, we establish bounds on the degree of regularity and the solving degree of a polynomial system that is regular in degree $D$. If the field size is smaller than the degree $D$ of the equations, then we establish bounds on the degree of regularity and the solving degree of the polynomial system to which we add the field equations. The bounds appear in Theorem 23, Corollary 27, Proposition 31, and Theorem 33.
@@ 2 Preliminaries
Let $\mathbb{F}$ be an arbitrary field and let $\mathbb{F}_{q}$ be the finite field of cardinality $q$. Let $R=\mathbb{F}[x_{1},\ldots,x_{n}]$ be the ring of polynomials in $n$ variables with coefficients in the field $\mathbb{F}$. Even though our main motivation comes from cryptography, most of the results in this paper do not depend on the ground field and will be stated and proven over an arbitrary field.
We start by recalling how a typical multivariate one-way function is constructed. Let $\mathbb{F}=\mathbb{F}_{q}$, $f_{1},\ldots,f_{m}\in R$, and consider the evaluation map
@@
To hide the structure of $\mathcal{F}$, one composes it with two random invertible linear maps $S:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n}$ and $T:\mathbb{F}_{q}^{m}\rightarrow\mathbb{F}_{q}^{m}$. One obtains $\mathcal{P}=T\circ\mathcal{F}\circ S$, a set of $m$ polynomials $p_{1},\ldots,p_{m}$ in $n$ variables over $\mathbb{F}_{q}$. The public key of the multivariate scheme is $\mathcal{P}=(p_{1},\ldots,p_{m})$ and the private key is $\{\mathcal{F},S,T\}$. The trapdoor consists of constructing $\mathcal{F}$ such that $\mathcal{F}^{-1}$ is efficiently computable. Notice that $\mathcal{P}$ should be hard to invert without the knowledge of $S,T$, in particular it should be hard to recover the structure of $\mathcal{F}$ from $\mathcal{P}$.
@@
Finding inverse images with respect to $\mathcal{P}$ corresponds to computing solutions of multivariate polynomial systems. It is well-known that this problem can be solved by computing a Gröbner basis of the system $\mathcal{P}$, we refer the interested reader to [16] for more detail. The first algorithm for computing Gröbner bases appeared in the doctoral thesis of Buchberger [13]. Modern algorithms for computing Gröbner bases are based on liner algebra and are often more efficient than Buchberger’s. Examples of linear-algebra-based algorithms are $F_{4}$ [28], $F_{5}$ [29], the XL Algorithm [23], MutantXL [14], and their variants. All these algorithms compute the reduced row echelon form of the Macaulay matrix of the system in a given degree, for one or more degrees.
@@
We now define Macaulay matrices, see also [8, Section 1.2]. Fix a term order on $R=\mathbb{F}[x_{1},\ldots,x_{n}]$ and let $\mathcal{F}=\{f_{1},\dots,f_{m}\}\subseteq R$ be a system of homogeneous polynomials. The columns of the homogeneous Macaulay matrix $M_{d}$ of $\mathcal{F}$ are labeled by the monomials of degree $d$ in $R$ and arranged in decreasing order. The rows of $M_{d}$ are labeled by polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})=d$. The entry of $M_{d}$ in position $(i,j)$ is the coefficient of the monomial of column $j$ in the polynomial corresponding to the $i$-th row.
@@
Let $f_{1},\dots,f_{m}$ be arbitrary polynomials (not necessarily homogeneous). The columns of the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}$ are labeled by the monomials of $R$ of degree $\leq d$, arranged in decreasing order. The rows of $M_{\leq d}$ correspond to polynomials of the form $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})\leq d$. The entries of $M_{\leq d}$ are defined as in the homogeneous case. The rationale behind the use of homogeneous Macaulay matrices for homogeneous systems is that, for a homogeneous system, the Macaulay matrix $M_{\leq d}$ is a block matrix with blocks $M_{d},\ldots,M_{0}$.
@@
In order to compute a Gröbner basis of the system $\mathcal{F}$, one computes the reduced row echelon form of the corresponding Macaulay matrix in one or more degrees. If the system is homogeneous, one uses homogeneous Macaulay matrices. For a sufficiently large degree, this produces a reduced Gröbner basis with respect to the chosen order. Some algorithms, such as MutantXL, use a variation called mutant strategy in the non-homogeneous case: If row reduction of the Macaulay matrix $M_{\leq d}$ produces new polynomials $g_{1},\ldots,g_{\ell}$ of degree strictly smaller than $d$, one appends to the reduction of $M_{\leq d}$ the polynomials $m_{i,j}g_{j}$, for every $j$ and every monomial $m_{i,j}\in R$ such that $\deg(m_{i,j}g_{j})\leq d$. Then one computes the reduced row echelon form again. Throughout the paper, we refer to algorithms that employ the mutant strategy as mutant algorithms and to the others as standard algorithms. For a description of the basic version of these algorithms, we refer to [16, Section 3.1] and [17, Section 1].
@@
The complexity of computing the reduced row echelon form of the Macaulay matrices $M_{\leq d}$ and $M_{d}$ depends on their size, hence on the degree $d$. This motivates the next definition.
@@
Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ and let $\tau$ be a term order on $R$. The solving degree of $\mathcal{F}$ with respect to $\tau$ is the least degree $d$ such that Gaussian elimination on the Macaulay matrix $M_{\leq d}$ produces a $\tau$-Gröbner basis of $\mathcal{F}$. We denote by $\operatorname{\mathrm{solv.deg}}_{\tau}^{s}(\mathcal{F})$ the solving degree of $\mathcal{F}$ with respect to a standard algorithm and by $\operatorname{\mathrm{solv.deg}}_{\tau}^{m}(\mathcal{F})$ the solving degree of $\mathcal{F}$ with respect to a mutant algorithm. When $\tau$ is the degree reverse lexicographic order, we omit the subscript $\tau$.
It is well-known that the solving degree is not invariant under coordinate change. In addition, the solving degree may depend on the algorithm used to perform the Gröbner basis computation. In particular, for mutant algorithms it may be smaller than for standard ones. Finally, the solving degree depends on the choice of a term order on $R$.
The complexity of linear-algebra-based algorithms is dominated by the cost of computing a degree reverse lexicographic Gröbner basis of the system, see [16, Sections 2 and 3] for more detail. Therefore, an upper bound on the solving degree with respect to the reverse lexicographic order yields an upper bound on the complexity of computing a lexicographic Gröbner basis, hence on the complexity of solving the polynomial system.
Let $I$ be a homogeneous ideal of $R$. For an integer $d\geq 0$, we denote by $R_{d}$ the vector space generated by the degree $d$ monomials of $R$. We denote by $I_{d}=I\cap R_{d}$ the $\mathbb{F}$-vector space of homogeneous polynomials of degree $d$ in $I$. For $g\in R$ a polynomial, we denote by $g^{\operatorname{\mathrm{top}}}$ the homogeneous part of $g$ of largest degree. For example, if $g=x^{3}+2xy^{2}-y+1\in\mathbb{F}[x,y]$, then $g^{\operatorname{\mathrm{top}}}=x^{3}+2xy^{2}$. For a polynomial system $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$, we denote by $\mathcal{F}^{\operatorname{\mathrm{top}}}\subseteq R$ the homogeneous system $\{f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}\}$. Up to performing Gaussian elimination on a matrix whose rows correspond to $f_{1},\ldots,f_{m}$, we may suppose that
$f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent.
Throughout the paper, we always make this assumption.
@@
In the cryptographic literature, the degree of regularity is often assumed to be finite and is used as a proxy for the solving degree. This is the case, e.g., in the GeMSS specification documents [19]. However, this does not always produce reliable estimates, even in the case when the degree of regularity is finite. In fact, there are examples in which the degree of regularity is much smaller than the solving degree, see e.g. [12, Examples 3.2 and 3.3]. However, a result by Semaev and Tenti [39, 41] shows that, under suitable assumptions, the solving degree of a system which contains the field equations is at most twice the degree of regularity. Due to this result, an upper bound for the degree of regularity yields a proven upper bound for the solving degree of standard algorithms.
Theorem 3 ([41, Corollary 3.67] and [39, Theorem 2.1]).
Let $\mathcal{F}=\{f_{1},\ldots,f_{m},x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\subseteq R$ be a polynomial system. If $d_{\operatorname{reg}}(\mathcal{F})\geq\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}$, then
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F})\leq 2d_{\operatorname{reg}}(\mathcal{F})-2.$
Notice that in most systems of cryptographic interest, the field size and the degrees of the polynomials are relatively small. Therefore, one expects that Theorem 3 applies to such systems.
@@
Theorem 4 ([38, Proposition 3.10]).
Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system, then
$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F})\leq\max\{d_{\operatorname{reg}}(\mathcal{F})+1,\deg(f_{1}),\ldots,\deg(f_{m})\}.$
Notice that, for both standard and mutant algorithms, it makes sense to assume that
$\operatorname{\mathrm{solv.deg}}(\mathcal{F})\geq\max\{\deg(f_{1}),\ldots,\deg(f_{m})\}.$ (1)
In fact, any equation of $\mathcal{F}$ of degree greater than the solving degree may be removed from the system without altering the Gröbner basis computation. This is of course delicate in practice, as one may not have an a priori estimate on the solving degree that allows them to eliminate equations from the system. Notice moreover that most multivariate systems associated to cryptographic schemes have low degree, so (1) is usually satisfied.
Another algebraic invariant connected to the solving degree of a polynomial system is the Castelnuovo-Mumford regularity. Throughout the paper, we denote by $\operatorname{reg}(I)$ the Castelnuovo-Mumford regularity of the ideal $I$. We refer the reader to [16, Section 3.4] and [17] for its definition and a discussion of its relationship with the other invariants. Notice that, if $d_{\operatorname{reg}}(\mathcal{F})<+\infty$, then
@@
3 Systems that contain a regular sequence in a given degree
Consider a system $\mathcal{F}$ of $m$ equations in $n$ variables. As we discussed in the previous section, whenever the degree of regularity of $\mathcal{F}$ is finite, an upper bound on it directly translates into an upper bound on the solving degree of the system, hence on the complexity of computing its solutions. In this paper, we prove upper bounds on the degree of regularity of a system $\mathcal{F}$ such that the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}$ contains a regular sequence of polynomials of degree $D$. In the rest of this section, we discuss the property of containing a regular sequence in a given degree and in which sense this is a generic property. The next definition formalizes the correspondence between homogeneous systems and points in a product of projective spaces.